# Unlocking State-Tracking in Linear RNNs Through Negative Eigenvalues

### v1

@@ B.2 Proof of Theorem 2
One Layer Let $\widehat{\bm{H}}_{k}$ and $\hat{y}_{k}:=\mathrm{cast}(\mathrm{dec}(\widehat{\bm{H}}_{k},x_{k}))$ be the finite precision versions of the state ${\bm{H}}_{k}$ and (scalar) output of a one-layer LRNN on the input ${\bm{x}}=x_{1}\dots x_{k}=1^{k}$. Let also $y_{k}=\mathbf{1}\{k\mod m=0\}$ be the correct output recognizing the word ${\bm{x}}$. We will show that if the assumptions on the eigenvalues are not satisfied, i.e. if for any $x$, every eigenvalue $\lambda$ of ${\bm{A}}(x)$ is either real or such that $|\lambda|<1$, then there exist $\overline{{\bm{H}}}_{1},\overline{{\bm{H}}}_{2}\in\mathbb{C}^{n\times n}$, $\bar{y}_{1},\bar{y}_{2}\in\mathbb{R}^{p}$ and $\tau\in\mathbb{N}$ such that for all $k\geq\tau$
@@
To prove (6), we can proceed similarly to Theorem 1. Indeed, using the $k$-th power formula for the Jordan Decomposition of the matrix ${\bm{A}}(1)$ with eigenvalues $\lambda_{1},\dots,\lambda_{s}$ we can prove that if $|\lambda_{i}|<1$ or $\lambda_{i}\in\mathbb{R}$ and $\lambda_{i}\geq 1$, then when $k\to\infty$ each element of the corresponding Jordan block of ${\bm{A}}(1)^{k}$ either converges to a single value or diverges to +$\infty$+ ∞. If instead $\lambda_{i}\in\mathbb{R}$ and $\lambda_{i}\leq-1$, the diagonal element of the corresponding Jordan block takes the form $c_{k}=(-1)^{k}|\lambda_{i}|^{k}$, while the ones above the diagonal take the form $z_{k}=\binom{k}{j}(-1)^{k-t}|\lambda_{i}|^{k-t}$ with $t,j\leq n$. It follows that if we let $\bar{c}\in\{1,\infty\}$, then
$\lim_{k\to\infty}c_{2k}$ $=\bar{c},\quad\lim_{k\to\infty}c_{2k+1}=-\bar{c},\quad\lim_{k\to\infty}z_{2k}=\infty,\quad\lim_{k\to\infty}z_{2k+1}=-\infty.$
Therefore we can apply the same reasoning of Theorem 1 using the finite precision assumption to show that there exist $\bar{\tau}\in\mathbb{N}$, $\overline{\bm{C}}_{1},\overline{\bm{C}}_{2},\overline{\bm{D}}_{1},\overline{\bm{D}}_{2}\in\mathbb{C}^{n\times d}$ such that for every $k\geq\tau$ we have
@@
Finally if for simplicity we consider $\tau\bmod 2=0$, we have that for $2k\geq\tau$
@@
where we note that the limit for $k\to\infty$ of the term inside $\mathrm{cast}$ is well defined. Thus there exist $\overline{\bm{H}}_{1},\overline{\bm{H}}_{2}\in\mathbb{C}^{n\times d}$ and $\bar{k}\geq\tau$ such that (6) is satisfied, concluding the proof for the case of a single layer.
@@
We notice that the sequences ${\bm{H}}_{2k}^{(2)}$ and ${\bm{H}}_{2k+1}^{(2)}$ are in a form similar to ${\bm{H}}_{k}$ of the first layer. If the assumption on the eigenvalues of the state-transition matrices of the second layer does not hold, this means that for all ${\bm{x}},{\bm{y}}$ then each eigenvalue of ${\bm{A}}^{(2)}({\bm{x}}){\bm{A}}^{(2)}({\bm{y}})$, including ${\bm{C}}_{1}$, is either real (but possibly negative) or has modulus strictly smaller than one. Therefore, we can proceed similarly to the case of one layer, i.e. using the powers of the Jordan canonical form of ${\bm{C}}_{1}$, to show that if we let $\widehat{\bm{H}}_{2k}^{(2)}$ and $\widehat{\bm{H}}_{2k+1}^{(2)}$ being the finite precision counterparts of ${\bm{H}}_{2k}^{(2)}$ and ${\bm{H}}_{2k+1}^{(2)}$, then there exist $\overline{\bm{H}}^{(2)}_{1},\overline{\bm{H}}^{(2)}_{2},\overline{\bm{H}}^{(2)}_{3},\overline{\bm{H}}^{(2)}_{4}\in\mathbb{C}^{n\times d}$, $\bar{k}_{2}\geq 0$ such that for every $k\geq\bar{k}$
@@
Therefore, for $k\geq\bar{k}_{2}$, the function $k\mapsto\overline{\bm{H}}^{(2)}_{k}$ will be periodic with period a divisor of four and hence no matter the choice of $\mathrm{dec}^{(2)}$, also the function $k\mapsto\hat{\bm{y}}^{(2)}_{k}$ will be periodic with period a divisor of $4$. Consequently, with two layers one can recognize the language $(1^{m})^{*}$ only when $m=1$, $m=2$, or $m=4$, since those are the only cases where $k\mapsto y_{k}$ has a period which is a divisor of $4$. Thanks to the assumption on the eigenvalues of the products of state-transition matrices, we can extend this argument inductively to the case of an LRNN with $L$ layers. In particular, for the $i$-th layer, the induction hypothesis is that we assume $k\mapsto{\bm{x}}_{k}^{(i)}$, mapping $k$ to the $k$-th input to the layer, to be periodic with period a divisor of $2^{i-1}$ for $k$ large enough. Hence, there will be $2^{i-1}$ subsequences of states containing powers of the product of $2^{i-1}$ state-transition matrices. From our hypothesis on the eigenvalues of products of state-transition matrices, such product will have only real eigenvalues and hence each subsequence will have 2 converging subsequences resulting in $k\mapsto{\bm{H}}_{k}^{(i)}$ and consequently $k\mapsto\hat{\bm{y}}_{k}^{(i)}$ and $k\mapsto{\bm{x}}_{k}^{(i)}$, for $k$ large enough, being periodic with period a divisor of $2^{i}$. Therefore, for the $L$-th layer, there exists $\bar{k}_{L}\geq 0$ such that for every $k\geq\bar{k}_{L}$, if we let $\hat{\bm{y}}^{(L)}_{k}$ be the output of the last layer, the function $k\mapsto\hat{\bm{y}}^{(L)}_{k}$ is periodic with a period which is a divisor of $2^{L}$ and thus it can recognize the language $(1^{m})^{*}$ only when $02^{L}\mod m=$ 0, which happens only when there exists $p\leq L$ such that $m=2^{p}$ and hence $m$ is a power of two, ending the proof. ∎
@@ C.1 Proof of Proposition 1
Third item Let ${\bm{N}}={\bm{C}}_{1}{\bm{C}}_{2}\cdots{\bm{C}}_{k}\in\mathcal{M}_{k}^{n}((-1,1])$, with ${\bm{C}}_{i}={\bm{I}}-\beta_{i}{\bm{k}}_{i}{\bm{k}}_{i}^{\top}$ with $\left\lVert k_{i}\right\rVert=1$ and $\beta_{i}\in[0,2)$. If ${\bm{N}}={\bm{I}}$ the statement is satisfied, otherwise, let $\mathcal{V}=\mathrm{span}\{{\bm{k}}_{i}\,:\,i\in\{1,\dots,k\},\beta_{i}>0\}$. Any unit vector ${\bm{v}}\in\mathbb{R}^{n}$ can then be written as ${\bm{v}}={\bm{v}}_{1}+{\bm{v}}_{2}$ with ${\bm{v}}_{1}\in\mathcal{V}$, ${\bm{v}}_{2}\in\mathcal{V}^{\top}$ and $\left\lVert{\bm{v}}_{1}\right\rVert,\left\lVert{\bm{v}}_{2}\right\rVert\leq 1$. Now, if ${\bm{v}}_{1}=0$, then ${\bm{N}}{\bm{v}}={\bm{v}}$, and hence ${\bm{v}}$ is an eigenvector with eigenvalue $1$. Instead, if ${\bm{v}}_{1}\neq 0$, then there exists $i^{\prime}\in\{1,\dots,k\}$ (we take the largest one one) such that $\beta_{i^{\prime}}\in(0,2)$ and ${\bm{v}}^{\top}{\bm{k}}_{i^{\prime}}={\bm{v}}_{1}^{\top}{\bm{k}}_{i^{\prime}}\in(0,1]$ and if $i^{\prime}<k$, then either $\beta_{j}=0$ or ${\bm{k}}_{j}^{\top}{\bm{v}}=0$ so that ${\bm{C}}_{j}{\bm{v}}={\bm{v}}$ for all $j\in\{i^{\prime}+1,\dots,k\}$. Moreover, we have that
$\lVert{\bm{C}}_{i^{\prime}}{\bm{v}}\rVert^{2}=\lVert{\bm{v}}-\beta_{i^{\prime}}{\bm{k}}_{i^{\prime}}{\bm{k}}_{i^{\prime}}^{\top}{\bm{v}}\rVert^{2}=1-\beta_{i^{\prime}}(2-\beta_{i^{\prime}})({\bm{v}}^{\top}{\bm{k}}_{i})^{2}<1,$
@@
$\left\lVert{\bm{N}}{\bm{v}}\right\rVert=\left\lVert{\bm{C}}_{1}{\bm{C}}_{2}\dots\bm{C}_{k}{\bm{v}}\right\rVert=\left\lVert{\bm{C}}_{1}{\bm{C}}_{2}\dots\bm{C}_{i^{\prime}}{\bm{v}}\right\rVert\leq\left\lVert{\bm{C}}_{1}\right\rVert\cdots\left\lVert{\bm{C}}_{i}{\bm{v}}\right\rVert<1.$
Therefore, if ${\bm{v}}$ is also an eigenvector with eigenvalue $\lambda\in\mathbb{C}$, then $\left\lVert{\bm{N}}{\bm{v}}\right\rVert=|\lambda|<1$. Hence, we proved that for every eigenvector with eigenvalue $\lambda$ either $\lambda=1$ or $|\lambda|<1$. It remains to show that all eigenvalues of ${\bm{N}}\in\mathcal{M}_{k}^{n}([0,1])$ are real. For $k=1$ it follows due to ${\bm{N}}$ being symmetric, for $k\geq 2$ let ${\bm{D}}={\bm{C}}_{1}{\bm{C}}_{2}\cdots{\bm{C}}_{k-1}$ so that ${\bm{N}}={\bm{D}}{\bm{C}}_{k}$ and let ${\bm{v}}$ be any eigenvector of ${\bm{N}}$ with eigenvalue $\lambda$ and $\left\lVert{\bm{v}}\right\rVert=1$. Then it holds that
${\bm{v}}^{\top}{\bm{C}}_{k}{\bm{N}}{\bm{v}}=\lambda{\bm{v}}^{\top}{\bm{C}}_{k}{\bm{v}}.$
Therefore if ${\bm{v}}^{\top}{\bm{C}}_{k}{\bm{v}}\neq 0$, then $\lambda={\bm{v}}^{\top}{\bm{C}}_{k}{\bm{N}}{\bm{v}}/{\bm{v}}^{\top}{\bm{C}}_{k}{\bm{v}}\in\mathbb{R}$. Otherwise when ${\bm{v}}^{\top}{\bm{C}}_{k}{\bm{v}}=0$ it follows that
${\bm{v}}^{\top}{\bm{C}}_{k}{\bm{v}}=\left\lVert{\bm{v}}\right\rVert^{2}-\beta_{k}({\bm{k}}_{k}^{\top}{\bm{v}})^{2}=1-\beta_{k}({\bm{k}}_{k}^{\top}{\bm{v}})^{2}=0,$
which is true only if $\beta_{k}=1$ and either ${\bm{v}}={\bm{k}}_{k}$ or ${\bm{v}}=-{\bm{k}}_{k}$ and thus ${\bm{C}}_{k}{\bm{v}}=\pm{\bm{C}}_{k}{\bm{k}}_{k}=0$ and hence $\lambda=0$, which concludes the proof. ∎
@@ C.2 Proof of Theorem 3
We view the LRNN layer in (1) as the automaton $\mathcal{A}_{\mathrm{lin}}=(\Sigma,\mathcal{H},{\bm{H}}_{0},\delta_{\mathrm{lin}})$, where $\delta_{\mathrm{lin}}({\bm{H}},w)={\bm{A}}(w){\bm{H}}+{\bm{B}}(w)$, which is extended in the usual way, and =$\mathcal{H}=\{\delta_{\mathrm{lin}}({\bm{H}}_{0},{\bm{w}})\,:\,{\bm{w}}\in\Sigma^{*}\}$. Since $\mathcal{T}(\mathcal{A})$ is a group, from Cayley’s theorem we have that it is isomorphic to a subgroup of $S_{n}$, which is the set of permutations on a set of $n$ elements. Furthermore, each element in $S_{n}$ can be represented as an $n\times n$ permutation matrix. Since in general $n\neq|Q|$, we cannot let $\mathcal{H}$ to be a set of one hot vectors each corresponding to states in $Q$. Instead, we let ${\bm{H}}_{0}=(1,\dots,n)^{\top}$, $\mathcal{P}\subset\{0,1\}^{n\times n}$ be the set of permutation matrices and set ${\bm{B}}\equiv 0$ and ${\bm{A}}:\Sigma\to\mathcal{P}$ to be the function mapping each letter $w\in\Sigma$ to the permutation matrix corresponding to $\delta(\cdot,w)$. With this choice we can see that the function $f:\mathcal{T}(\mathcal{A}_{\mathrm{lin}})\to\mathcal{T}(\mathcal{A})$ such that $f(\delta_{\mathrm{lin}}(\cdot,{\bm{w}}))=\delta(\cdot,{\bm{w}})$ for every ${\bm{w}}\in\Sigma^{*}$ is one-to-one (biejctive), and from our choice of ${\bm{H}}_{0}$, the map $h:\mathcal{T}(\mathcal{A}_{\mathrm{lin}})\to\mathcal{H}$ such that for every ${\bm{w}}\in\Sigma^{*}$, $h(\delta_{\mathrm{lin}}(\cdot,{\bm{w}}))=\delta_{\mathrm{lin}}({\bm{H}}_{0},{\bm{w}})$ is also bijective. Moreover, the map $\phi:\mathcal{T}(\mathcal{A})\to Q$ such that $\phi(\delta(\cdot,{\bm{w}}))=\delta(q_{0},{\bm{w}})$ is surjective because we consider states that are only reachable from the initial state $q_{0}$, i.e. $Q=\{\delta(q_{0},{\bm{w}})\,:\,{\bm{w}}\in\Sigma^{*}\}$. Hence if we set $g=\phi\circ f\circ h^{-1}$, then $g:\mathcal{H}\to Q$ is surjective and for every $w\in\Sigma$ and ${\bm{H}}\in\mathcal{H}$ we have that
@@
Thus, we have shown that such LRNN implements $\mathcal{A}$ and it does so with finite precision because the entries of all vectors and matrices are bounded integers. Moreover, Let $k=\max_{w\in\Sigma}\sum_{q\in Q}\mathbf{1}\{\delta(q,w)\neq q\}=\max_{w\in\Sigma}\sum_{i=1}^{n}\mathbf{1}\{({\bm{A}}(w){\bm{H}}_{0})_{i}={\bm{H}}_{0,i}\}$ be the maximum number of displaced element of the permutation associated with the alphabet $\Sigma$. Then, this means that each permutation can be written as a product of at most $k-1$ permutations of two elements. Hence, for every $w\in\Sigma$, ${\bm{A}}(w)\in\mathcal{M}_{k-1}^{n}(\{-1,1\})$.
If in addition there exists $m\in\mathbb{N}$ such that $\mathcal{T}(\mathcal{A})$ is isomorphic to a subgroup of the cyclic group $\mathbb{Z}_{m}$ with elements $\{0,\dots,m-1\}$, we can modify the construction above to use a smaller dimension. If $m=2$, then $\mathbb{Z}_{2}$ has elements $\{0,1\}$, and $\mathcal{A}$ implements the parity automaton. Thus, we can set ${\bm{H}}_{0}=-1$, ${\bm{A}}(0)=1$, ${\bm{A}}(1)=-1$ and $g(1)=1$ while $g(0)=1$, which means that we can use a scalar recursion. Otherwise, if $m\geq 3$, we can modify the construction above by setting ${\bm{H}}_{0}=(1,0)^{\top}$ and, if for simplicity we assume $\Sigma\in\{0,\dots,m-1\}$, for every $w\in\Sigma$ we let ${\bm{A}}(w)$ be the $2\times 2$ rotation matrix corresponding to $\delta(\cdot,w)$:

### v2

@@ B.2 Proof of Theorem 2
One Layer Let $\widehat{\bm{H}}_{k}$ and $\hat{y}_{k}:=\mathrm{cast}(\mathrm{dec}(\widehat{\bm{H}}_{k},x_{k}))$ be the finite precision versions of the state ${\bm{H}}_{k}$ and (scalar) output of a one-layer LRNN on the input ${\bm{x}}=x_{1}\dots x_{k}=1^{k}$. Let also $y_{k}=\mathbf{1}\{k\mod m=0\}$ be the correct output recognizing the word ${\bm{x}}$. We will show that if the assumptions on the eigenvalues are not satisfied, i.e. if for any $x$, every eigenvalue $\lambda$ of ${\bm{A}}(x)$ is real, then there exist $\overline{{\bm{H}}}_{1},\overline{{\bm{H}}}_{2}\in\mathbb{C}^{n\times n}$, $\bar{y}_{1},\bar{y}_{2}\in\mathbb{R}^{p}$ and $\tau\in\mathbb{N}$ such that for all $k\geq\tau$
@@
To prove (7), we set
$\widehat{\bm{H}}_{k}=\mathrm{cast}\left(\sum_{i=0}^{k-1}\mathrm{cast}\left({\bm{A}}(1)^{i}{\bm{B}}(1)\right)+\mathrm{cast}\left({\bm{A}}(1)^{k}{\bm{H}}_{0}\right)\right),$
and proceed similarly to Theorem 1. Indeed, using the $k$-th power formula for the Jordan Decomposition of the matrix ${\bm{A}}(1)$ with eigenvalues $\lambda_{1},\dots,\lambda_{s}$, the imaginary and real part of each element of the matrices ${\bm{A}}(1)^{k}{\bm{B}}(1)$ and ${\bm{A}}(1)^{k}{\bm{H}}_{0}$ will be a linear combination of elements of the Jordan blocks taking the same form of $a_{k}$ in Lemma 1. Therefore since our assumptions with $L=1$ imply that $\lambda_{i}\in\mathbb{R}$ for every $i$, we can apply Lemma 1 to show that there exist $\bar{\tau}\in\mathbb{N}$, $\overline{\bm{C}}_{1},\overline{\bm{C}}_{2},\overline{\bm{D}}_{1},\overline{\bm{D}}_{2}\in\mathbb{C}^{n\times d}$ such that for every $k\geq\tau$ we have
@@
Finally, if for simplicity we consider $\tau\bmod 2=0$, we have that for $2k\geq\tau$
@@
where by factoring out $k$ inside $\mathrm{cast}$, we note that for large enough $k$, the real and imaginary parts of each element of the matrices inside $\mathrm{cast}$ will be either constant, smaller than $\min_{x\in\mathbb{R}}\mathrm{cast}(x)$ or larger than $\max_{x\in\mathbb{R}}\mathrm{cast}(x)$. Thus there exist $\overline{\bm{H}}_{1},\overline{\bm{H}}_{2}\in\mathbb{C}^{n\times d}$ and $\bar{k}\geq\tau$ such that (7) is satisfied, concluding the proof for the case of a single layer.
@@
We notice that the sequences ${\bm{H}}_{2k}^{(2)}$ and ${\bm{H}}_{2k+1}^{(2)}$ are in a form similar to ${\bm{H}}_{k}$ of the first layer. If the assumption on the eigenvalues of the state-transition matrices of the second layer does not hold, this means that for all ${\bm{x}},{\bm{y}}$ each eigenvalue of ${\bm{A}}^{(2)}({\bm{x}}){\bm{A}}^{(2)}({\bm{y}})$, including ${\bm{C}}_{1}$, is real (but possibly negative). Therefore, we can proceed similarly to the case of one layer, i.e. using the powers of the Jordan canonical form of ${\bm{C}}_{1}$, to show that if we let $\widehat{\bm{H}}_{2k}^{(2)}$ and $\widehat{\bm{H}}_{2k+1}^{(2)}$ being the finite precision counterparts of ${\bm{H}}_{2k}^{(2)}$ and ${\bm{H}}_{2k+1}^{(2)}$, then there exist $\overline{\bm{H}}^{(2)}_{1},\overline{\bm{H}}^{(2)}_{2},\overline{\bm{H}}^{(2)}_{3},\overline{\bm{H}}^{(2)}_{4}\in\mathbb{C}^{n\times d}$, $\bar{k}_{2}\geq 0$ such that for every $k\geq\bar{k}$
@@
Therefore, for $k\geq\bar{k}_{2}$, the function $k\mapsto\overline{\bm{H}}^{(2)}_{k}$ will be periodic with period a divisor of four and hence no matter the choice of $\mathrm{dec}^{(2)}$, also the function $k\mapsto\hat{\bm{y}}^{(2)}_{k}$ will be periodic with period a divisor of $4$. Consequently, with two layers one can recognize the language $(1^{m})^{*}$ only when $m=1$, $m=2$, or $m=4$, since those are the only cases where $k\mapsto y_{k}$ has a period which is a divisor of $4$. Thanks to the assumption on the eigenvalues of the products of state-transition matrices, we can extend this argument inductively to the case of an LRNN with $L$ layers. In particular, for the $i$-th layer, the induction hypothesis is that we assume $k\mapsto{\bm{x}}_{k}^{(i)}$, mapping $k$ to the $k$-th input to the layer, to be periodic with period a divisor of $2^{i-1}$ for $k$ large enough. Hence, there will be $2^{i-1}$ subsequences of states containing powers of the product of $2^{i-1}$ state-transition matrices. From our hypothesis on the eigenvalues of products of state-transition matrices, such product will have only real eigenvalues and hence each subsequence will have 2 converging subsequences resulting in $k\mapsto{\bm{H}}_{k}^{(i)}$ and consequently $k\mapsto\hat{\bm{y}}_{k}^{(i)}$ and hence $k\mapsto{\bm{x}}_{k}^{(i+1)}$, for $k$ large enough, being periodic with period a divisor of $2^{i}$. Therefore, for the $L$-th layer, there exists $\bar{k}_{L}\geq 0$ such that for every $k\geq\bar{k}_{L}$, the function $k\mapsto\hat{\bm{y}}^{(L)}_{k}$ is periodic with a period which is a divisor of $2^{L}$ and thus it can recognize the language $(1^{m})^{*}$ only when $02^{L}\mod m=$ 0, which happens only when there exists $p\leq L$ such that $m=2^{p}$ and hence $m$ is a power of two, ending the proof. ∎
@@ C.1 Proof of Proposition 1
Third item Let ${\bm{N}}={\bm{C}}_{1}{\bm{C}}_{2}\cdots{\bm{C}}_{k}\in\mathcal{M}_{k}^{n}((-1,1])$, with ${\bm{C}}_{i}={\bm{I}}-\beta_{i}{\bm{z}}_{i}{\bm{z}}_{i}^{\top}$ with $\left\lVert{\bm{z}}_{i}\right\rVert=1$ and $\beta_{i}\in[0,2)$. If ${\bm{N}}={\bm{I}}$ the statement is satisfied, otherwise, let $\mathcal{V}=\mathrm{span}\{{\bm{z}}_{i}\,:\,i\in\{1,\dots,k\},\beta_{i}>0\}$. Any unit vector ${\bm{v}}\in\mathbb{R}^{n}$ can then be written as ${\bm{v}}={\bm{v}}_{1}+{\bm{v}}_{2}$ with ${\bm{v}}_{1}\in\mathcal{V}$, ${\bm{v}}_{2}\in\mathcal{V}^{\top}$ and $\left\lVert{\bm{v}}_{1}\right\rVert,\left\lVert{\bm{v}}_{2}\right\rVert\leq 1$. Now, if ${\bm{v}}_{1}=0$, then ${\bm{N}}{\bm{v}}={\bm{v}}$, and hence ${\bm{v}}$ is an eigenvector with eigenvalue $1$. Instead, if ${\bm{v}}_{1}\neq 0$, then there exists $i^{\prime}\in\{1,\dots,k\}$ (we take the largest one one) such that $\beta_{i^{\prime}}\in(0,2)$ and ${\bm{v}}^{\top}{\bm{z}}_{i^{\prime}}={\bm{v}}_{1}^{\top}{\bm{z}}_{i^{\prime}}\in(0,1]$ and if $i^{\prime}<k$, then either $\beta_{j}=0$ or ${\bm{z}}_{j}^{\top}{\bm{v}}=0$ so that ${\bm{C}}_{j}{\bm{v}}={\bm{v}}$ for all $j\in\{i^{\prime}+1,\dots,k\}$. Moreover, we have that
$\lVert{\bm{C}}_{i^{\prime}}{\bm{v}}\rVert^{2}=\lVert{\bm{v}}-\beta_{i^{\prime}}{\bm{z}}_{i^{\prime}}{\bm{z}}_{i^{\prime}}^{\top}{\bm{v}}\rVert^{2}=1-\beta_{i^{\prime}}(2-\beta_{i^{\prime}})({\bm{v}}^{\top}{\bm{z}}_{i^{\prime}})^{2}<1,$
@@
$\left\lVert{\bm{N}}{\bm{v}}\right\rVert=\left\lVert{\bm{C}}_{1}{\bm{C}}_{2}\dots\bm{C}_{k}{\bm{v}}\right\rVert=\left\lVert{\bm{C}}_{1}{\bm{C}}_{2}\dots\bm{C}_{i^{\prime}}{\bm{v}}\right\rVert\leq\left\lVert{\bm{C}}_{1}\right\rVert\cdots\left\lVert{\bm{C}}_{i^{\prime}}{\bm{v}}\right\rVert<1.$
Therefore, if ${\bm{v}}$ is also an eigenvector with eigenvalue $\lambda\in\mathbb{C}$, then $\left\lVert{\bm{N}}{\bm{v}}\right\rVert=|\lambda|<1$. Hence, we proved that for every eigenvector with eigenvalue $\lambda$ either $\lambda=1$ or $|\lambda|<1$.
It remains to show that all eigenvalues of ${\bm{N}}\in\mathcal{M}_{2}^{n}([0,1])$ are in $[0,1]$. From the assumptions ${\bm{N}}={\bm{C}}_{1}{\bm{C}}_{2}$ with ${\bm{C}}_{1},{\bm{C}}_{2}$ symmetric and positive semi-definite, therefore ${\bm{C}}_{1}$ has a unique symmetric and positive semi-definite square root ${\bm{C}}_{1}^{1/2}$ such that ${\bm{C}}_{1}^{1/2}{\bm{C}}_{1}^{1/2}={\bm{C}}_{1}$. If ${\bm{C}}_{1}$ is non-singular (invertible) then
${\bm{C}}_{1}{\bm{C}}_{2}={\bm{C}}_{1}^{1/2}{\bm{C}}_{1}^{1/2}{\bm{C}}_{2}{\bm{C}}_{1}^{1/2}{\bm{C}}_{1}^{-1/2}.$
Thus, ${\bm{C}}_{1}{\bm{C}}_{2}$ is similar to ${\bm{C}}_{1}^{1/2}{\bm{C}}_{2}{\bm{C}}_{1}^{1/2}$ and shares its eigenvalues. Moreover ${\bm{C}}_{1}^{1/2}{\bm{C}}_{2}{\bm{C}}_{1}^{1/2}$ is symmetric positive semi-definite (having real nonnegative eigenvalues) because ${\bm{C}}_{1}^{1/2}$ and ${\bm{C}}_{2}$ are symmetric and ${\bm{v}}^{\top}{\bm{C}}_{1}^{1/2}{\bm{C}}_{2}{\bm{C}}_{1}^{1/2}{\bm{v}}={\bm{z}}^{\top}{\bm{C}}_{2}{\bm{z}}\geq 0$ with ${\bm{z}}={\bm{C}}_{1}^{1/2}{\bm{v}}$ since ${\bm{C}}_{2}$ is positive semi-definite. Instead, if ${\bm{C}}_{1}$ is singular, for $t>0$ the matrix ${\bm{C}}_{1}+t{\bm{I}}$ is positive definite and non-singular. Hence $({\bm{C}}_{1}+t{\bm{I}}){\bm{C}}_{2}$ has real and nonnegative eigenvalues. Since ${\bm{C}}_{1}{\bm{C}}_{2}=\lim_{t\to 0}({\bm{C}}_{1}+t{\bm{I}})({\bm{C}}_{2})$ and the eigenvalues are a continuous function of the entry of the matrix, ${\bm{C}}_{1}{\bm{C}}_{2}$ has positive real eigenvalues. Since the modulus of any eigenvalue is smaller or equal than the euclidean norm of the matrix, which is smaller than one from the first point of the theorem, the statement follows. ∎
@@ C.2 Proof of Theorem 3
We view the LRNN layer in (1) as the automaton $\mathcal{A}_{\mathrm{lin}}=(\Sigma,\mathcal{H},{\bm{H}}_{0},\delta_{\mathrm{lin}})$, where $\delta_{\mathrm{lin}}({\bm{H}},w)={\bm{A}}(w){\bm{H}}+{\bm{B}}(w)$, which is extended in the usual way, and =$\mathcal{H}=\{\delta_{\mathrm{lin}}({\bm{H}}_{0},{\bm{w}})\,:\,{\bm{w}}\in\Sigma^{*}\}$. Since we assumed that $\mathcal{T}(\mathcal{A})$ is a group, from Cayley’s theorem we have that it is isomorphic to a subgroup of $S_{n}$, which is the set of permutations on a set of $n$ elements. Furthermore, each element in $S_{n}$ can be represented as an $n\times n$ permutation matrix. Since in general $n\neq|Q|$, we cannot let $\mathcal{H}$ to be a set of one hot vectors each corresponding to states in $Q$. Instead, we let ${\bm{H}}_{0}=(1,\dots,n)^{\top}$, $\mathcal{P}\subset\{0,1\}^{n\times n}$ be the set of permutation matrices and set ${\bm{B}}\equiv 0$ and ${\bm{A}}:\Sigma\to\mathcal{P}$ to be the function mapping each letter $w\in\Sigma$ to the permutation matrix corresponding to $\delta(\cdot,w)$. With this choice we can see that the function $f:\mathcal{T}(\mathcal{A}_{\mathrm{lin}})\to\mathcal{T}(\mathcal{A})$ such that $f(\delta_{\mathrm{lin}}(\cdot,{\bm{w}}))=\delta(\cdot,{\bm{w}})$ for every ${\bm{w}}\in\Sigma^{*}$ is one-to-one (biejctive), and from our choice of ${\bm{H}}_{0}$, the map $h:\mathcal{T}(\mathcal{A}_{\mathrm{lin}})\to\mathcal{H}$ such that for every ${\bm{w}}\in\Sigma^{*}$, $h(\delta_{\mathrm{lin}}(\cdot,{\bm{w}}))=\delta_{\mathrm{lin}}({\bm{H}}_{0},{\bm{w}})$ is also bijective. Moreover, the map $\phi:\mathcal{T}(\mathcal{A})\to Q$ such that $\phi(\delta(\cdot,{\bm{w}}))=\delta(q_{0},{\bm{w}})$ is surjective because we consider states that are only reachable from the initial state $q_{0}$, i.e. $Q=\{\delta(q_{0},{\bm{w}})\,:\,{\bm{w}}\in\Sigma^{*}\}$. Hence if we set $g=\phi\circ f\circ h^{-1}$, then $g:\mathcal{H}\to Q$ is surjective and for every $w\in\Sigma$ and ${\bm{H}}\in\mathcal{H}$ we have that
@@
Thus, we have shown that such an LRNN implements $\mathcal{A}$ and it does so with finite precision because the entries of all vectors and matrices are bounded integers. Moreover, Let $k=\max_{w\in\Sigma}\sum_{q\in Q}\mathbf{1}\{\delta(q,w)\neq q\}=\max_{w\in\Sigma}\sum_{i=1}^{n}\mathbf{1}\{({\bm{A}}(w){\bm{H}}_{0})_{i}={\bm{H}}_{0,i}\}$ be the maximum number of displaced element of the permutation associated with the alphabet $\Sigma$. Then, this means that each permutation can be written as a product of at most $k-1$ permutations of two elements. Hence, for every $w\in\Sigma$, ${\bm{A}}(w)\in\mathcal{M}_{k-1}^{n}(\{-1,1\})$.
If in addition there exists $m\in\mathbb{N}$ such that $\mathcal{T}(\mathcal{A})$ is isomorphic to a subgroup of the cyclic group $\mathbb{Z}_{m}$ with elements $\{0,\dots,m-1\}$, we can modify the construction above to use a smaller dimension. If $m=2$, then $\mathbb{Z}_{2}$ has elements $\{0,1\}$, and $\mathcal{A}$ implements the parity automaton. Thus, we can set ${\bm{H}}_{0}=-1$, ${\bm{A}}(0)=1$, ${\bm{A}}(1)=-1$ and $g(1)=1$ while $g(-1)=0$, which means that we can use a scalar recursion. Otherwise, if $m\geq 3$, we can modify the construction above by setting ${\bm{H}}_{0}=(1,0)^{\top}$ and, if for simplicity we assume $\Sigma\in\{0,\dots,m-1\}$, for every $w\in\Sigma$ we let ${\bm{A}}(w)$ be the $2\times 2$ rotation matrix corresponding to $\delta(\cdot,w)$: